OpenStudy (anonymous):

given f(x)=x^2+7 and g(x)=(x-4)/x find (g*f)(-1)

OpenStudy (freckles):

plug in -1 into both expressions evaluate then multiply the results

OpenStudy (anonymous):

f(-1)=-1^2+7 g(-1)=(-1-4)/-1?

OpenStudy (freckles):

f(-1)=(-1)^2+7 g(-1)=(-1-4)/(-1) so I made one correct to f(-1) expression thing anyways you need to evaluate both (-1)^2+7 and (-1-4)/(-1)

OpenStudy (freckles):

and then multiply the results

OpenStudy (anonymous):

6 and 5?

OpenStudy (anonymous):

so is it 30?

OpenStudy (irishboy123):

sometimes \[( f * g)(x) \equiv f (g(x)) \]

OpenStudy (anonymous):

so I got 29=-30